OpenStudy (anonymous):

What is 2x+7/10=3/5?

TheSmartOne (thesmartone):

Welcome to Open Study!!

TheSmartOne (thesmartone):

Is your question \[\frac{ 2x+7 }{ 10}=\frac{ 3 }{ 5}\]

TheSmartOne (thesmartone):

@sashafeirce